A)An object of volume 7.50×10−4m3 and density 1.15×103kg/m3 is completely submerged in a fluid. Calculate the weight of this object.

B)Calculate the buoyant force on this object if the fluid is air (density 1.20 kg/m3).

C)Calculate the buoyant force on this object if the fluid is water (density 1.00×103kg/m3).

D)Calculate the buoyant force on this object if the fluid is glycerin (density 1.26×103kg/m3)
Answer #1

A] We know that weight W = mg = rho*V*g = 1.15e3*7.5e-4*9.8 = 8.4525 N

B] Buoyant force = rhoair*Vg = 1.2*7.5e-4*9.8 = 0.00882 N

C] Buoyant force = rhowater*Vg = 1e3*7.5e-4*9.8 = 7.35 N

D] Buoyant force = rhoglycerin*Vg = 1.26e3*7.5e-4*9.8 = 9.261 N